# Concentration of Measure for Distributions Generated via Diffusion Models

### v2

@@ 3.4 Limitations of current universality results
Assumptions 1 hold, for example, for any sub-Gaussian $\mathbf{x}$ with mean and covariance satisfying $\|\bm{\mu}\|_{2}=O(1)$ and $\frac{c\mathbf{I}_{d}}{d}\leq\bm{\Sigma}^{\frac{1}{2}}\leq\frac{C\mathbf{I}_{d}}{d}$ (see Remark 5 in Ghane et al. [2024] for details). However, assuming that $\frac{c\mathbf{I}_{d}}{d}\leq\bm{\Sigma}^{\frac{1}{2}}\leq\frac{C\mathbf{I}_{d}}{d}$ is crucial here, as otherwise one can take a Gaussian $\mathbf{x}$ with $\bm{\Sigma}=diag(1,\frac{1}{4},\dots,\frac{1}{d^{2}})$ and $\bm{\mu}=0$ and notice that $Var(\|\mathbf{x}\|^{2})=\text{Tr}(\bm{\Sigma}^{2})-\text{Tr}(\bm{\Sigma})^{2}$ converges to a strictly positive number, violating the third part of Assumptions 1 for $\mathbf{C}=\mathbf{I}_{d}$, while $\mathbf{x}$ is normalized correctly in the sense that $\mathbb{E}_{\mathbf{x}}\|\mathbf{x}\|^{2}=\text{Tr}(\bm{\Sigma})=O(1)$.
@@
Unfortunately, as can be seen in Figures 9 and 10, the spectra of diffusion-generated images look qualitatively similar to the "power law" $\bm{\Sigma}=diag(1,\frac{1}{4},\dots,\frac{1}{d^{2}})$, meaning that Theorem 1 does not apply in this setting. Moreover, to the best of the authors’ knowledge, such covariance matrices break the assumptions commonly made in papers focusing on universality for regression, which is usually simpler to study. For example, Montanari and Saeed [2022] also have to assume $\frac{c\mathbf{I}_{d}}{d}\leq\bm{\Sigma}^{\frac{1}{2}}\leq\frac{C\mathbf{I}_{d}}{d}$ to get concrete results for over-parametrized regression (cf. Theorem 5 in Montanari and Saeed [2022]). Despite this, as can be seen in the next section, the universality of the classification error does not break thus posing an interesting challenge of relaxing Assumptions 1 in Theorem 1.
While, technically speaking, universality is proven only for the objectives of the form (3), in practice one usually adds a softmax function $S(\mathbf{z}_{1},\dots,\mathbf{z}_{k})=(\dots,\frac{e^{\mathbf{z}_{\ell}}}{\sum e^{\mathbf{z}_{i}}},\dots)$
$\min_{\mathbf{W}\in\mathbb{R}^{d\times k}}\|\mathbf{W}-\mathbf{W}_{0}\|_{F}^{2}$ (5)
$s.t\quad S(\mathbf{X}\mathbf{W})\approx\mathbf{Y}$
@@
Here, the approximate equality comes from the fact that the coordinates of the range of the softmax cannot turn exactly into zeros but will be very close to it on the training data if one fits the objective (5). Since this objective is of much greater practical interest than (3) and has better convergence properties, we add softmax into the objective for numerical validation of universality in the next section. Note that, from theoretical standpoint, it raises the question of incorporating softmax into the framework of Theorem 1.
3.5 Main Results
To explain why diffusion models do not perform well at generating heavy-tailed data, we prove the following result:
@@ 3.5 Main Results
Assume that the denoiser $D_{\theta}(\mathbf{x}^{(i)},t_{i:i+1})$ is trained in such a way that $\|\mathcal{R}^{(i)}_{D_{\theta}}(\mathbf{x}^{(i)},t_{i:i+1})\|_{2}\leq\|\mathbf{x}^{(i)}\|_{2}$ under the notation from (2) holds for every sampling step with high probability w.r.t the randomness in $\bm{\epsilon}_{i}$. Then the resulting output $\mathbf{x}^{(N)}$ satisfies the CoM property from Definition 1 for $C=2$ and $\sigma=t_{0}$, where $\mathbf{x}^{(0)}\sim\mathcal{N}(0,t_{0}^{2}\mathbf{I}_{d})$.
@@
See Appendix B. ∎
The assumption $\|\mathcal{R}_{D_{\theta}}^{(i)}(\mathbf{x},t_{i:i+1})\|_{2}\leq\|\mathbf{x}\|_{2}$ might come out as very specific. In addition, it was not clear to us how to analytically verify that it is true. Nevertheless, we justify it by the following empirical observation. Understanding mathematically why Empirical Observation 1 holds thus poses an interesting challenge as well.
Empirical Observation 1.
Each sampling step $\mathbf{x}^{(i+1)}=\mathcal{R}^{(i)}_{D_{\theta}}(\mathbf{x}^{(i)},t_{i:i+1})$ of the Algorithm 1 decreases norms, i.e. $\|\mathbf{x}^{(i)}\|\leq\|\mathbf{x}^{(i-1)}\|$ is satisfied throughout the reverse process. The results of the corresponding experiments can be found in Figure 4.
We observe this contractivity in the sampling process for the setting described in Section 3.2. This observation raises the possibility that many diffusion models used in practice may also possess a contractive sampling process. An important future direction is to understand the conditions under which this property holds, given the base and target distributions, noise schedule, and denoiser training.
As explained in Subsection 3.3, CoM property is insufficient for concluding universality from any of the known universality theorems unless the upper bound from the right hand side of Definition 1 goes to 00. Nevertheless, our experiments suggest that universality holds for diffusion-generated images despite this technicality. As such, we would like to report it as an empirical observation and present the question of extending Theorem 1 to capture more complicated covariance matrices $\bm{\Sigma}$ such as the power law as an open question for future theory works.
@@
The distributions of images generated via EDM diffusion models satisfy Gaussian Universality of the test error in the sense of the conclusion of Theorem 1 for weights trained via minimizing $\|\mathbf{Y}-S(\mathbf{X}\mathbf{W})\|_{2}^{2}$ using SGD. The experiments are presented in Figure 6 preceded by the description of the setup.

### v3

@@ 3.2 Limitations of current universality results
Assumptions 1 hold, for example, for any sub-Gaussian $\mathbf{x}$ with mean and covariance satisfying $\|\bm{\mu}\|_{2}=O(1)$ and $\frac{c\mathbf{I}_{d}}{d}\leq\bm{\Sigma}\leq\frac{C\mathbf{I}_{d}}{d}$ (see Remark 5 in Ghane et al. (2024) for details). However, assuming that $\frac{c\mathbf{I}_{d}}{d}\leq\bm{\Sigma}\leq\frac{C\mathbf{I}_{d}}{d}$ is crucial here, as otherwise one can take a Gaussian $\mathbf{x}$ with $\bm{\Sigma}=diag(1,\frac{1}{4},\dots,\frac{1}{d^{2}})$ and $\bm{\mu}=0$ and notice that $Var(\|\mathbf{x}\|_{2}^{2})=\text{Tr}(\bm{\Sigma}^{2})-\text{Tr}(\bm{\Sigma})^{2}$ converges to a strictly positive number, violating the third part of Assumptions 1 for $\mathbf{C}=\mathbf{I}_{d}$, while $\mathbf{x}$ is normalized correctly in the sense that $\mathbb{E}_{\mathbf{x}}\|\mathbf{x}\|^{2}=\text{Tr}(\bm{\Sigma})=O(1)$.
@@
Unfortunately, as can be seen in Figure 4, the spectra of diffusion-generated images look qualitatively similar to the ”power law” $\bm{\Sigma}=diag(1,\frac{1}{4},\dots,\frac{1}{d^{2}})$, meaning that Theorem 1 does not apply in this setting. Moreover, to the best of the authors’ knowledge, such covariance matrices break the assumptions commonly made in papers focusing on universality for regression, which is usually simpler to study. For example, Montanari and Saeed (2022) also have to assume $\frac{c\mathbf{I}_{d}}{d}\leq\bm{\Sigma}\leq\frac{C\mathbf{I}_{d}}{d}$ to get concrete results for over-parametrized regression (cf. Theorem 5 in Montanari and Saeed (2022)). Despite this, in the next section, the experimental results seem to suggest the universality of the classification error does not break, This motivates us to relax the Assumptions 1 in Theorem 1. While, technically speaking, universality is proven only for objectives of the form (1), in practice one usually adds a softmax function $S(\mathbf{z}_{1},\dots,\mathbf{z}_{k})=(\dots,\frac{e^{\mathbf{z}_{\ell}}}{\sum e^{\mathbf{z}_{i}}},\dots)$
$\min_{\mathbf{W}\in\mathbb{R}^{d\times k}}\|\mathbf{W}-\mathbf{W}_{0}\|_{F}^{2}\quad s.t\quad S(\mathbf{X}\mathbf{W})\approx\mathbf{Y}$
@@
Here, the approximate equality comes from the fact that the coordinates of the range of the softmax cannot turn exactly into zeros but will be very close to it on the training data if one fits the objective (3.2). Since this objective is of much greater practical interest than (1) and has better convergence properties, we add softmax into the objective for numerical validation of universality in the next section. Note that, from theoretical standpoint, it raises the question of incorporating softmax into the framework of Theorem 1.
Empirical Observation 1.
The test error of the weights trained via minimizing $\|\mathbf{Y}-S(\mathbf{X}\mathbf{W})\|_{F}^{2}$ on images generated via EDM diffusion models matches the test error of the weights trained on the matching GMM. The experiments are presented in Figure 1 preceded by the description of the setup.
3.3 Approximate Concentration of Measure Property
In this section, we provide additional insights leading to Empirical Observation 1. Before doing so, we present a definition central to the result of this section. We use the following definition of concentration. Informally, it means that the tails of the distribution decay exponentially fast. Note that it corresponds to the Lipschitz Concentration Property for $q=2$ from Seddik et al. (2020), by setting $c=0$.
Definition 1 (ACoM).
Given a probability distribution $x\sim\mathbb{P}$ where $\mathbf{x}\in\mathbb{R}^{d}$, we say that $\mathbf{x}$ satisfies the Approximate Concentration of Measure Property (ACoM) if there exist absolute constants $\Bigl(C,c,c^{\prime},d,\sigma\Bigr)>0$ such that for any $L$-Lipschitz function $f:\mathbb{R}^{d}\to\mathbb{R}$ and $s>0$ it holds that
$\mathbb{P}\Bigl(|f(\mathbf{x})-\mathbb{E}f(\mathbf{x})|>s\Bigr)\leq Ce^{-(\frac{s}{L\sigma})^{2}}+ce^{-c^{\prime}d}$ (2)
The distributions satisfying ACoM arise naturally in many applications and are quite ubiquitous. We appeal to the following proposition proven in Rudelson and Vershynin (2013) :
Proposition 1.
The distribution $\mathbf{x}\sim\mathcal{N}(0,\bm{\Sigma})$ satisfies the ACoM property 1. Moreover, the corresponding $C=2$, $c=0$, and $\sigma=\|\bm{\Sigma}^{\frac{1}{2}}\|_{op}$.
Note that by taking $c=0$, we recover the classical notion of Concentration of Measure in the literature. If $\bm{\Sigma}=\frac{\mathbf{I}_{d}}{d}$ and $f(\mathbf{x})=\|\mathbf{x}\|_{2}$, then Proposition 1 implies the classical fact that the norm of the normalized standard vector converges to $1$ in probability as $d\to\infty$ because in this case the upper bound of Definition 2 becomes $2e^{-(\frac{s}{\sigma})^{2}}=2e^{-sd}\to 0$. However, if $\bm{\Sigma}$ is also normalized as $\text{Tr}(\bm{\Sigma})=1$ but $\|\bm{\Sigma}\|_{op}=\Theta(1)$ (which happens, for instance, if the ordered eigenvalues of $\bm{\Sigma}$ follow the power law $\lambda_{i}=\tilde{C}i^{-\alpha}$ for some $\tilde{C}>0$ and $\alpha>1$), then the variance of $\mathbf{x}$ does not have to go to zero anymore, but Definition 1 still implies that $\mathbf{x}$ has exponentially decaying tails (to be more precise, $\mathbf{x}$ is a sub-Gaussian random vector–see Definition 3.4.1 in Vershynin (2018)). Gaussians are far from the only distributions satisfying ACoM; other examples include the strongly log-concave distributions, and the Haar measure-we refer to Section 5.2 in Vershynin (2018) for more examples. The concentration of measure phenomenon has played a key role in the development of many areas such as random functional analysis, compressed sensing, and information theory.
Diffusion: We denote the denoiser used in the forward and backward processes of a conventional diffusion model as $D_{\theta}(\cdot)$. In this paper, we focus only on the sampling process where we use $\mathbf{x}^{(i)}$ to denote the sample generated at the $i$’the step of the sampling process starting from the initialization $\mathbf{x}^{(0)}\sim\mathcal{N}(0,t_{0}^{2}\mathbf{I}_{d})$. We run the sampling for $N$ steps. We use $t_{i}$ to denote the value of the noise scale used in the sampling process. Furthermore, $\mathbf{x}^{(i)}\mapsto\mathcal{R}^{(i)}_{D_{\theta}}(\mathbf{x}^{(i)},t_{i:i+1})=\mathbf{x}^{(i+1)}$ is used to denote the mapping that generates the sample in the next step and we use $t_{i:i+1}$ to summarize $(t_{i},t_{i+1})$. Building on Definition 1, the following result provides insights on why the test error of generalized linear models trained on diffusion-generated images matches that of the matching GMM.
@@ 3.3 Approximate Concentration of Measure Property
(a) Assume that $\|\nabla_{\mathbf{x}}D_{\theta}(\mathbf{x}^{(i)})\|\leq L_{D}$ holds with probability at least $1-c_{1}e^{-c_{2}d}$ for some $L_{D}>0$ and all $i=0,\dots,N$. Then there exists $L_{\mathcal{R}}>0$, such that the following holds as well:
$\mathbb{P}\Bigl(\|\nabla_{\mathbf{x}}\mathcal{R}^{(i)}_{D_{\theta}}(\mathbf{x}^{(i)},t_{i:i+1})\|_{2}\leq L_{\mathcal{R}}\Bigr)\geq 1-c_{1}e^{-c_{2}d}$
(b) Furthermore, if $L_{\mathcal{R}}\leq 1$, then the resulting output $\mathbf{x}^{(N)}$ satisfies the following tail bound for every $L_{f}$-Lipschitz $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and every $s>0$:
$\mathbb{P}\Bigl(|f(\mathbf{x}^{(N)})-\mathbb{E}$ $f(\mathbf{x}^{(N)})|>s\Bigr)\leq 2\exp\Big(-\frac{s^{2}}{2L_{f}^{2}}\Bigr)+2Nc_{1}e^{-c_{2}d}$
@@
See Appendix C. ∎
Remark 1.
Theorem 2 implies that diffusion generated distributions satisfy the ACoM property from Definition 1 as long as $N=O(e^{c_{3}d})$ for some $c_{3}<c_{2}$. This makes our results applicable to the scenario where the number of steps in the sampling procedure is exponential in the dimension of the data. To motivate the assumption $L_{\mathcal{R}}\leq 1$ made in part (b) of Theorem 2, we would like to refer the reader to Figure 3 as a partial empirical justification for it. We would also like to point out that the norms decrease in such a manner that, in fact, $\|\mathbf{x}^{(i)}-\tilde{\mathbf{x}}^{(i)}\|_{2}\ll\|\mathbf{x}^{(0)}-\tilde{\mathbf{x}}^{(0)}\|_{2}\>$ for two independent noise initializations $\mathbf{x}^{(0)},\tilde{\mathbf{x}}^{(0)}\sim\mathcal{N}(0,t_{0}^{2}\mathbf{I}_{d})$ and the corresponding points $\mathbf{x}^{(i)},\tilde{\mathbf{x}}^{(i)}$ obtained from $\mathbf{x}^{(0)},\tilde{\mathbf{x}}^{(0)}$ by applying $i$ consecutive steps of the backward process.
We have also made the following empirical observation, that partly supports the assumption $L_{\mathcal{R}}\leq 1$ from Theorem 2 as well. Understanding mathematically why Empirical Observation 2 holds thus poses an interesting challenge.
@@
Each sampling step $\mathbf{x}^{(i+1)}=\mathcal{R}^{(i)}_{D_{\theta}}(\mathbf{x}^{(i)},t_{i:i+1})$ of the Algorithm 1 in Appendix A decreases norms, i.e. $\|\mathbf{x}^{(i)}\|_{2}\leq\|\mathbf{x}^{(i-1)}\|_{2}$ throughout the reverse process. The results of the corresponding experiments can be found in Figure 3.
We observe this contractivity in the sampling process for the setting described in Appendix A. This observation raises the possibility that many diffusion models used in practice may also possess a contractive sampling process. An important future direction is to understand the conditions under which this property holds, given the base and data distributions, noise schedule, and denoiser training.
We finally note that the ACoM property in (2) is insufficient for concluding universality from any of the known universality theorems even for $c=0$ unless the upper bound from the right-hand side of Definition 1 goes to 0 as $d\rightarrow\infty$. Nevertheless, our experiments suggest that universality holds for diffusion-generated images despite this technicality. As such, we report it as an empirical observation and present the question of extending Theorem 1 to capture more complicated covariance matrices $\bm{\Sigma}$ such as the power law as an open question for future theory works.